OpenStudy (anonymous):

Simplify 1+cos(x+pi/2)+cos(x-pi/2)

OpenStudy (anonymous):

1 + 2 cos x

OpenStudy (anonymous):

Steps?

OpenStudy (anonymous):

cos(a+b)=cos a cos b -sin a sin b understand what I mean ?

OpenStudy (anonymous):

Yes .

OpenStudy (anonymous):

But if cos(pi/2) is 0 , then how come the answer is 1+2cosx?

OpenStudy (anonymous):

hey erkoT

OpenStudy (anonymous):

?

OpenStudy (anonymous):

want to help me out with three questions??

OpenStudy (anonymous):

If I have any idea on how to solve them , then sure .

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). Answer y = one third x + 7 y = -3x + 3 y = 'one third x + 3 y = -3x + 7

OpenStudy (anonymous):

y = -3x + 3

OpenStudy (anonymous):

if so ecaplin how u did it please

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). Answer y = 2x + 6 y = -one half x + 3 y = -one half x + 6 y = 2x + 3

OpenStudy (anonymous):

So basically the first thing you want to do is to find the ascent of the line , which is -3 from the formula , y=kx+b (Not sure how it's in your country) but from here you'll get that k or the ascent is -3 , Then use the formula y-y'=k(x-x') , so you'll get y-6=-3(x+1) -> y=-3x+3 . Sorry for my typing speed .

OpenStudy (anonymous):

oh its okay and thank u actually i get it for once.

OpenStudy (anonymous):

i got one more after the one up there! Thats if u dont mind.. and where are you from?

OpenStudy (anonymous):

y = -one half x + 6 , want me to explain too?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

And i'm from Estonia .

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). Answer y = four thirds x + 8 y = four thirds x + 3 y = -three fourths x + 8 y = -three fourths x + 3

OpenStudy (anonymous):

y = four thirds x + 8

OpenStudy (anonymous):

where is that...europe?

OpenStudy (anonymous):

Yes , you're correct .

OpenStudy (anonymous):

and do u mins checking on for me?

OpenStudy (anonymous):

Checking on?

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, -1). Answer y = one third x - 3 y = one third x + 17 y = -3x - 3 y = -3x + 17 this is a?

OpenStudy (anonymous):

im pretty sure that this one is a ????

OpenStudy (anonymous):

A

OpenStudy (anonymous):

You're correct . Well done .

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

You're welcome .

OpenStudy (anonymous):

i have more but its a little different and im sure ur tired and i know u helped me enough already!

OpenStudy (anonymous):

i became ur fan!

OpenStudy (anonymous):

Thank you ! We could also become friends .

OpenStudy (anonymous):

?

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

But send me one of your problems you have left . I'll see how different is it . Well , Facebook for example .

OpenStudy (anonymous):

ugh i dont have a fb

OpenStudy (anonymous):

Which of the following is the slope between the points (0, -10)and (-4, 0)?

OpenStudy (anonymous):

Im warring you that i have like 40 questions...

OpenStudy (anonymous):

left to be answered and u can say no

OpenStudy (anonymous):

? u there

OpenStudy (anonymous):

The slope or ascent as I told it before if calculateable with the formula . k=(y2-y1)/(x2-x1) . So just plug in the values k=(0-(-10))/(-4-0) , which is -2,5 .

OpenStudy (anonymous):

Which of the following is the slope between the points (3, -2)and (8, 4)? Answer -five sixths -six fifths six fifths five sixths

OpenStudy (anonymous):

Yea , i'm still here . But i'm afraid I can't help you any more today cause I need to go . >But I'll help you wiht the last one .

OpenStudy (anonymous):

ok and thanks.

OpenStudy (anonymous):

k=(4-(-2))/(8-3)=6/5

OpenStudy (anonymous):

thanks so much what time is it there???

OpenStudy (anonymous):

13.04 or 1pm (not sure about the pm) .

OpenStudy (anonymous):

LOL THANKS AGAIN byee (:

OpenStudy (anonymous):

See you ! :)